$19.30 13.90 83.80 S 26.80 18.90 105.00 S 2,170 6,300 20 120 1,500 5003.000
2. С Unit Cost 3. per unit for each product under the tradtional costing system. (Round your Intermediate calculations and final answers to 2 decimal places.) he activity rates d Keller wanted to implement an ABC system. (Round your answers to 2
your Cost Cost
media%2Fda4%2Fda445d9a-fa98-4e2f-968e-35
Answer #1
Sandy Beach Rocky River
Direct Material per unit 19.3 26.8
Direct Labor per unit 13.9 18.9
Selling Price per unit 83.8 105
Expected production per month 1220 units 920 units
Monthly overhead Total 11560 Cost Driver Sandy Beach Rocky River Total
Setup costs 2170 No. of Setups 20 15 35
Quality control 6300 No. of inspections 120 405 525
Maintenance 3090 No. of machine hours 1500 1500 3000
1
Under traditional costing, all overheads are allocated on the basis of machine hours. i.e. 1500 and 1500.
Allocated overhead costs = Total overheads/ Total machine hours
Allocated overheads costs = 11560/3000 = $3.85
Overheads allocated = 3.85*1500each = $ 5780
Sandy Beach Rocky River
Overheads as per traditional method 5780 5780
2
Production cost per unit = (Direct labor per unit + Direct Material per unit + (Overheads/Production units)
Sandy Beach Rocky River
Production cost per unit 13.9+19.3+(5780/1220) 18.9+26.8+(5780/920)
Production cost per unit $ 37.93 Approx $ 51.98 Approx
3
Gross margin per unit = Selling price per unit - production costs per unit
Gross Margin % = Gross Margin/ Selling Price*100
Sandy Beach Rocky River
Gross Margin 83.8-37.94 = 45.86 105-51.98 = 53.02
Gross Margin% 54.72% 50.49%
4
Cost Pool Cost Driver
Setup No. of setups
Quality Control No. of Inspections
Maintenance No. of machine hours
Setup costs = Total setup costs / Total setup * No. of setup of each activity
Qualty control costs = Total Quality control costs/Total inspections * No. of inspections for each activity
Maintenance costs = Total maintenance costs/Total machine hours * No. of machine hours for each activity
Sandy Beach Rocky River
Setup 2170/35*20 = 1240 2170/35*15 = 930
Qualtiy Control 6300/525*120 = 1440 6300/525*405 = 4860
Maintenance 3090/3000*1500 = 1545 3090/3000*1500 = 1545
5
Taking demand as the cost driver for overheads
Total Demand = 1220 + 920 = 2140
Sandy Beach Rocky River
Setup 2170/2140*1220 = 1237 2170/2140*920 = 933
Qualtiy Control 6300/2140*1220 = 3592 6300/2140*920 = 2708
Maintenance 3090/2140*1220 = 1762 3090/2140*920 = 1328
Total Overhead Costs 6591 4969
6
Production cost per unit = Direct Material + Direct Labor + Setup Cost + Quality Control + Maintenance
Costs of cost pools have already been calculated by ABC in Question 4
Sandy Beach Rocky River
Direct Material 19.3 26.8
Direct Labor 13.9 18.9
Setup 1240/1220 = 1.01 930/920 = 1.01
Quality Control 1440/1220 = 1.18 4860/920 = 5.28
Maintenance 1545/1220 = 1.27 1545/920 = 1.68
Total cost per unit 36.67 53.67
7
Gross margin per product = Selling price per unit - cost per unit
Gross margin 5 = Selling price per unit - Cost per unit/Selling price per unit * 100
Sandy Beach Rocky River
Gross Margin 83.8-36.67 = 47.13 105-53.67 = 51.33
Gross Margin% 56.24% 48.88%
8
Sandy Beach Rocky River
Gross Margin per unit (Traditional) 45.86 53.02
Gross margin per unit (ABC) 47.13 51.33
Calculated in question 3,7 above